OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - 2x - 3? x = -3, 1 x = -3, -1 x = 3, 1 x = 3, -1

OpenStudy (anonymous):

@Lady.Liv1776 @kiamousekia @Blonde_Gangsta

OpenStudy (anonymous):

@foreverjennnicole

OpenStudy (anonymous):

first, do you have any ideas on how you think you should solve this, taco?

OpenStudy (anonymous):

do u think id be here if i did enchilada

OpenStudy (anonymous):

Use the caret ^ shift 6 to raise to the powers f(x) = x^3 - x^2 - 2x............factor out"x" the GCF f(x) = x(x^2 - x - 2).............what numbers multily to = 2 and add = -1 f(x) = x (x - 2)(x + 1) x = 0......x - 2= 0........x + 1= 0 x = 0..........x = 2.............x = - 1....the three zeros of f(x)

OpenStudy (anonymous):

so is it d

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

I may not be correct though